# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

### v1

@@ 5.3 Effectiveness of AQUA
InternVL3 Naive $2.86$ $27.11\%$ $0.41$ $4.56$ $17.12\%$ $0.65$
Opt. $1.45$ $19.34\%$ $5.39e^{-3}$ $1.90$ $19.45\%$ $3.87e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $75.72\%$ $1.49e^{-50}$ $1.85$ $72.46\%$ $4.31e^{-26}$
Qwen-VL -Chat Naive $2.86$ $15.79\%$ $0.59$ $4.56$ $5.71\%$ $0.91$
Opt. $1.45$ $21.29\%$ $9.05e^{-3}$ $1.90$ $18.91\%$ $1.21e^{-3}$
[…]
Qwen2.5- VL-Instruct Naive $2.86$ $38.15\%$ $0.25$ $4.56$ $15.87\%$ $0.86$
Opt. $1.45$ $19.96\%$ $7.35e^{-3}$ $1.90$ $18.51\%$ $6.77e^{-3}$
[…]
$\textit{AQUA}_{spatial}$ $1.29$ $98.42\%$ $8.29e^{-72}$ $1.85$ $89.85\%$ $2.92e^{-49}$
In this section, we evaluate the effectiveness of our AQUA. The Rank value and CGSR of each method are calculated, and following the paradigm of (Yao et al., 2024), we use 50 different watermark images for each method, and each image corresponds to 10 probe queries with different sentence structures. The experiment is repeated 10 times to obtain the p-values in the Table 1. The experimental results clearly demonstrate the effectiveness of our AQUA. Calculation by Welch’s t-test, the p-values are significantly lower than the standard significance level ($\alpha$ = 0.05). This allows us to confidently reject the null hypothesis $\mathcal{H}_{0}:\mu_{suspect}=\mu_{clean}$, providing strong statistical evidence that the AQUA method enables reliable detection of the injected watermarks.
Figure 4: TPR vs. FPR of two methods and two baselines.
Figure 5: The relationship between p-value and query times.
p-value vs. Query times. While all methods can eventually produce statistically significant results (i.e., low p-values) to reject the null hypothesis, the number of queries required to reach significance is a crucial factor in real-world applications—especially when each query may incur cost or be subject to limitations. To evaluate query efficiency, we measure how efficiently each method achieves statistical significance. Figure 5 shows that $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ can obtain a p-value less than the significance level within 30 queries, while the Opt. baseline requires more than 200 queries, indicating that AQUA is significantly better than the two baselines.
FPR vs. TPR. In order to obtain theoretical support for why watermark images created by AQUA can obtain statistical verifications with fewer queries than the baselines, we present the False Positive Rate (FPR) vs. True Positive Rate (TPR) plot (Figure 5). FPRs are gained from measuring the generator’s (LLaVA-NeXT) output without watermark images in the database, and TPRs are obtained with 1, 2, 3, 5, and 10 watermark images for each probe query. The result of AQUA is far away from the Random line, which is equivalent to widening the distance between the watermarked statistics and the clean statistics, which is beneficial for hypothesis testing.
Real-world Deployment. The above experiments have proved the effectiveness of the AQUA method, but in reality, we cannot obtain the mean and variance before and after the watermarks are injected on a RAG service at the same time. We can only get one mean and variance ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from the suspected RAG service, so we propose a verification strategy with a predefined VSR’s reference distribution. We first characterize the reference distribution of a clean Multimodal RAG using mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and the same with a watermarked one ($\mu_{wm}$, $\sigma^{2}_{wm}$). Subsequently, we can perform Welch’s t-test between ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) and two respective reference distributions. The null hypotheses ($\mathcal{H}_{0}$) for two hypothesis tests are: Suspect vs. Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$ and Suspect vs. Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$. To avoid a false accusation, the significance level $\alpha$ can be set to a very low value (e.g. $3e^{-5}$ in (Jovanović et al., 2025)). Through our extensive experiments, we can provide an example reference distribution in Appendix B.
5.4 Harmlessness of AQUA
Normal Query. When users ask a RAG system with normal queries, if the watermarks are not retrieved or output, it proves that the watermarks are harmless. We use normal queries (e.g. “What animals race in the Kentucky Derby?”) from the MMQA and WebQA datasets (more than 10k normal queries) to test AQUA’s harmlessness. Experiments show that when using a normal query within just one watermark image in the knowledge base, the retrieval rate of both $\textit{AQUA}_{acronym}$’s and $\textit{AQUA}_{spatial}$’s watermark images are 0%, and the CGSR is also 0% on four generators. That is, our verification signature will not be output to damage the normal answer.
Table 2: Examples of relevant queries and corresponding results.
Type Probe Query Relevant Query Rank SimScore $\uparrow$
Acronym-
replace
[…]
Acronym-
no_instruction
What is the subtitle of UGP? What is UGP? $1.07$ $70.18\%$
Spatial-
imprecise
What fruit is the monkey holding like a phone? What is the monkey holding? $2.93$ $75.87\%$
Relevant Query. Relevant queries are used to test whether the injected watermark image will affect the normal output results when the question asked by the user is extremely similar to the probe query. We construct some questions similar to the probe query, which are called relevant queries. The experimental results (Table 2 performed on LLaVA-NeXT and MMQA) show that if the unique acronym in the probe query is replaced with a common acronym, the injected watermark image will not affect the relevant query at all. Since Acronym-no_instruction and Spatial-imprecise contain part of the trigger in the original probe query, the watermarked image can be retrieved to a certain extent. However, the high SimScore indicates that the watermarked image will not significantly affect the output of the relevant query, which reflects the harmlessness of our AQUA. More results are shown in Appendix C.
5.5 Stealthiness of AQUA
(a) Image embeddings
(b) Text embeddings
Figure 6: PCA comparison on embeddings. Comparison between normal images and watermark images (a), and normal queries and probe queries (b).
PCA Visualization. Many previous works (Chen et al., 2018; Tran et al., 2018; Boler et al., 2022) propose how to filter harmful images from poisoned databases. Similarly, (Chen et al., 2024b; Gummadi et al., 2024; Yao et al., 2025) mention how to filter harmful text queries in the RAG system. So, if the embeddings of watermark images and probe queries are similar to the embeddings of original images and normal queries in the dataset, detecting and filtering the watermark images is difficult for Adversary. We employ Principal Component Analysis (PCA) (Pearson, 1901; Hotelling, 1933) to visualize the embeddings of watermark images and their probe queries compared to those of original images and normal queries from the MMQA dataset (Fig. 6). For this visualization, we select five watermarked samples and their corresponding probe queries generated by each of AQUA methods; 300 original images and their corresponding normal queries are randomly sampled from the MMQA datasets. These results indicate that AQUA maintains strong stealthiness while preserving high retrieval performance.
Retrieval Ratio vs. Watermark Number. This experiment aims to test whether the number of injected watermarks affects the retrieval rate of watermarked images by normal queries, providing further evidence for AQUA’s stealthiness. Figure 7 shows that as the number of injected watermark images increases, the retrieval ratio of watermark images constructed by our $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ are both less than 0.1% and does not increase significantly, further demonstrating the stealthiness of our approach.
5.6 Robustness of AQUA
{textblock*}
0.63 (1pt,1pt)
Table 3: The Rank and p-value values of the watermark image after the following transformations.
Transformations $\textit{AQUA}_{acronym}$ $\textit{AQUA}_{spatial}$
Rank $\downarrow$ p-value $\downarrow$ Rank $\downarrow$ p-value $\downarrow$
Rescale $1.03$ $7.57e^{-274}$ $1.36$ $3.62e^{-62}$
Rotate $1.07$ $3.76e^{-142}$ $1.61$ $3.41e^{-45}$
Gaussian $1.07$ $4.10e^{-264}$ $1.46$ $1.64e^{-56}$
Rescale + Rotate + Gaussian $1.06$ $1.70e^{-256}$ $1.79$ $1.43e^{-39}$
We assess the robustness of our designed watermark images against potential image transformations, such as rescaling, rotating, and adding Gaussian noise. As in the setting of experiment 5.3, the p-value is obtained through 512 query trials. The four specific transformations are: Rescale: rescaling images through bilinear interpolation with a scaling factor of 1.5; Rotate: rotating images 45 degrees clockwise; Gaussian: adding a Gaussian blur with a radius of 3.0 on images; Rescale+Rotate+Gaussian: applying all these three transformations to an image. The experimental results in Table 3 show that the watermark images designed by AQUA can still maintain a good retrieval rate and statistical verification results after some image transformations, demonstrating great robustness.
6 Discussion
Limitation. Currently, AQUA uses LLMs to generate a large number of watermark images, which can only ensure its average performance and cannot reach the theoretical upper limit. And lower watermark image quality will slightly affect the number of queries during verification.
Future Works. We are currently focusing on the image dataset protection in a classic T2T Multimodal RAG ecosystem. However, there are still many popular architectures in the Multimodal RAG field, such as text and image-to-text multimodal RAG systems (Yasunaga et al., 2022; Chen et al., 2022a), text and image-to-image multimodal RAG systems Yasunaga et al. (2022); Chen et al. (2022b); Shalev-Arkushin et al. (2025), and so on. Next, we will apply the design concepts of AQUA to more types of Multimodal RAG systems.
@@ 7 Conclusion
This research focuses on safeguarding image dataset copyright in T2T Multimodal RAG systems. We proposed AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness. Two complementary watermarking strategies in AQUA can protect the copyright of image datasets through statistical verification methods using only a few watermark images. Since AQUA is the first method to protect data copyright through watermarking in the realistic black-box Multimodal RAG scenarios, we consider that AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
@@ A.1 Prompt
Acronym Prompt. Here is an example prompt (Figure 8) where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names. {num_of_watermark} indicates the number of watermarks you want.
Here is an example:
(UGP, Unicorn Grammar Parser)
Please create {num_of_watermark} pairs of uncommon acronyms and their
full names based on this example.
Figure 8: Prompt used to construct a large number of acronym pairs.
SimScore Prompt. This is the prompt (Figure 9) where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar. {Clean_Answer} and {Watermark_Answer} indicate the response of Multimodal RAG’s generator with the un-watermarked dataset and the watermarked one, respectively.
Determine the semantic similarity between the following two strings and
give your score on a scale of 0-100:
String 1: {Clean_Answer}
String 2: {Watermark_Answer}
Just answer with numbers.
Figure 9: Prompt used to obtain the SimScore between two strings.
@@ A.2 Watermark Examples
In this section, we will provide several classic watermark examples. The examples of $\textit{AQUA}_{acronym}$ method is in the Table 4 and $\textit{AQUA}_{spatial}$ in Table 6. The sentence structure and wording of the probe query can be freely selected, but semantic consistency must be ensured.
Table 4: Examples of watermark images designed by the $\textit{AQUA}_{acronym}$ method.
@@ 1. What’s the meaning of UGP?
1. Output the meaning of BLT.
2. What does the acronym BLT stand for based on the images?
3. Please output the subtitle of acronym BLT.
Bouncing Llama Technologies
1. Please state the full phrase that BLT abbreviates.
2. Could you provide the expansion for the acronym BLT?
3. What words do the letters in BLT represent?
Quantum Walrus Xylophone
1. What is the meaning of the acronym TPB?
2. Could you spell out what TPB represents?
3. Give me the full form of TPB.
Temporal Platypus Bagpipe
@@ 2. Could you decode XCO?
3. Give me the full form of TPB.
@@
Table 5: Examples of watermark images designed by the $\textit{AQUA}_{spatial}$ method.
@@ 1. Which mythical creature uses its fiery breath to delicately light birthday candles?
Appendix B Reference Distribution
$\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$ need to use different means and variances to characterize their respective reference distributions. Since this reference distribution is related to the specific watermark image constructed and its performance, here we can give an example reference distribution through our extensive experiments:
  • $\textit{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.005, 0.02); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.6, 0.2)
  • $\textit{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = (0.2, 0.2); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = (0.55, 0.25)
Appendix C More Results of Harmlessness of AQUA

### v2

@@ 5.6 Robustness of AQUA
System-Level Variations. In addition to image-level distortions and regeneration attacks, we further evaluate the stability of AQUA against practical deployment variations, such as benign system updates and corpus expansion. As detailed in Appendix C.4, altering the retrieval model (from CLIP to SigLIP (Zhai et al., 2023a)) or fine-tuning it yields only marginal changes to the retrieval Rank (Table 15). Furthermore, when the retrieval corpus is substantially enlarged by introducing 50k WebQA distractor images, the detection performance remains highly robust, achieving a CGSR of 96.76% for $\textbf{AQUA}_{acronym}$ and 91.85% for $\textbf{AQUA}_{spatial}$ (Table 16). These results confirm that AQUA maintains reliable performance under common real-world system modifications and data scaling.
Generator-Level Adaptations. We further consider adversaries who attempt to circumvent the watermark at the generator level. First, we test whether frontier VLMs (Gemini-3-Pro, GPT-5.1-High, Qwen3-VL(Google DeepMind, 2025; OpenAI, 2025; Bai and others, 2025b) refuse to process watermarked images as “unnatural” inputs; no refusal behavior is observed across all models. Second, we simulate an adversarial fine-tuning attack using LoRA (Hu et al., 2022) to train the generator to specifically reject watermark-related queries. While CGSR decreases moderately, the watermark remains reliably detectable, and the fine-tuning incurs a significant utility degradation ($8.62\%$ on MMQA accuracy), making it an impractical strategy for adversaries. Full details are provided in Appendix C.4.1.
@@ 6 Conclusion
This research focuses on safeguarding the copyright of image datasets in T2T Multimodal RAG systems. We propose AQUA, a watermarking framework that meets four design requirements: effectiveness, harmlessness, stealthiness, and robustness. The two complementary watermarking strategies in AQUA enable copyright protection for image datasets via statistical verification using only a few watermark images. Since AQUA is the first method to protect data copyright through watermarking in realistic black-box Multimodal RAG scenarios, AQUA can serve as a crucial baseline for future studies in Multimodal RAG data protection, contributing to more robust copyright protection in this important area.
Acknowledgments
We sincerely thank the anonymous reviewers for their detailed and constructive feedback. We are also grateful to the area chairs for their dedication in carefully handling the review process despite the significantly increased workload during ICLR 2026. This work is supported by Shanghai Engineering Research Center of Intelligent Vision and Imaging. We also thank members of the ASPIRE Lab at SIST, ShanghaiTech University, for helpful conversations during the course of this work.
@@ A.1 Prompt
Acronym Prompt Here is an example: (UGP, Unicorn Grammar Parser) Please create {num_of_watermark} pairs of uncommon acronyms and their full names based on this example.
This is an example prompt where we use LLM (Gemini-2.5-Pro) to build a large number of acronyms and their full names. {num_of_watermark} indicates the number of watermarks you want.
SimScore Prompt Determine the semantic similarity between the following two strings and give your score on a scale of 0-100: String 1: {Clean_Answer} String 2: {Watermark_Answer} Just answer with numbers.
This is the prompt where we ask LLM (Gemini-2.5-Pro) to determine whether two strings are semantically similar. {Clean_Answer} and {Watermark_Answer} indicate the response of Multimodal RAG’s generator with the un-watermarked dataset and the watermarked one, respectively.
VLM-based Filter Prompt for $\textbf{AQUA}_{acronym}$ You are an OCR filter. As long as the picture contains unusual acronym, it is a watermark picture. Your task is to identify if any of these five images are suspected of being the watermarked images. If yes, please answer ’yes’ and indicate how many watermarked images there are. If no, please answer ’no’.
VLM-based Filter Prompt for $\textbf{AQUA}_{spatial}$ You are a dataset watermark filter. Watermarked images are those that look unnatural or even defy common sense. You need to identify if any of these five images are suspected watermarks. If yes, please answer ’yes’ and indicate how many watermarked images there are. If no, please answer ’no’.
@@ A.2 Watermark Examples
The acronym images used by $\textbf{AQUA}_{acronym}$ can be generated through various means: they may be constructed manually, implemented programmatically via libraries, or produced with the assistance of Large Language Models (LLMs). There are no specific stylistic requirements for these images. They can range from simple black text on a white background, as exemplified above, to acronyms embedded within naturalistic images, as depicted in the Table 5. The sentence structure and wording of the probe query can be freely selected, but semantic consistency must be ensured.
Table 5: Examples of watermark images designed by the $\textbf{AQUA}_{acronym}$ method.
@@ 1. What’s the meaning of UGP?
1. UGP is the name of a store. What is the full name of this store?
2. Please output all the purple English letters on a black shop in a magic style
Unicorn Grammar Parser
1. Help me find a picture of a street with green plants. Please output the English words on the sign to the right of the green plants.
2. Find an image with a green UGP acronym and output the English text on the street sign in full.
Unicorn Grammar Parser
1. Find a green book standing on the table and output the title of the book.
2. What is the full name of a book titled UGP?
Unicorn Grammar Parser
@@ 2. Could you decode XCO?
3. Give me the full form of XCO.
@@
Here are three additional examples of the $\textbf{AQUA}_{spatial}$ method (Figure 7). Since the data provider adds the watermarks themselves, they have the flexibility to define watermarks that either slightly deviate from the dataset’s overall distribution or conform to it while containing subtle variations in detail. We provide a pipeline to guide data providers in achieving a balance between semantic distinctiveness and statistical naturalness for their specific datasets. The details of this pipeline are elaborated in Appendix A.4.
Table 6: Examples of watermark images designed by the $\textbf{AQUA}_{spatial}$ method.
@@ 1. Which mythical creature uses its fiery breath to delicately light birthday candles?
A.3 Visual Stealthiness Inspection
To further demonstrate the high stealthiness and visual naturalness of AQUA, we provide an extended gallery of watermarked samples in Figure 6. The gallery is organized into three columns: the first column displays images generated via $\textbf{AQUA}_{acronym}$, the second column shows $\textbf{AQUA}_{spatial}$, and the third column presents original, benign images from the MMQA dataset for comparison. As illustrated, the watermarked images maintain high visual fidelity. To the naked eye, they appear indistinguishable from the standard benign images, verifying that the watermarks do not introduce noticeable artifacts or unnatural distortions that would compromise the visual quality of the dataset.
AQUAacronym
AQUAspatial
MMQA (Normal)
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Extended Gallery of Watermarked vs. Normal Images. The figure presents 10 examples for each category, arranged in two columns per group. Columns 1-2: $\textbf{AQUA}_{acronym}$; Columns 3-4: $\textbf{AQUA}_{spatial}$; Columns 5-6: Original MMQA images. The high visual fidelity ensures indistinguishability from benign samples.
A.4 Image Caption Template and Evaluation Pipeline
We note that semantic distinctiveness in images is largely driven by object-level details, while naturalness is influenced by global factors such as composition, spatial arrangement, color, and background. Our image generation process is designed to preserve both aspects.
To ensure semantic distinctiveness, we employ structured and controllable templates. We then populate these templates with concept pairs that exhibit low co-occurrence probabilities—these can be identified using methods such as word embedding similarity, BERT-based measures, or prompting large language models.
Image Caption Template {number 1} {color 1} {object 1} {location/action/state} {number 2} {color 2} {object 2}
Any part of this template can be used as the verification signal, for instance, “what {object 2} is”, or “the {number 1} of object 1”. If using the example from our paper, the template would be “{a} {red} {apple} {on the head of} {a} {} {dog}”, and the verification signal would be “apple”.
To ensure statistical naturalness, we score the generated captions using pre-trained language models (e.g., BERT) and filter out those with high perplexity, which typically correspond to unnatural or ungrammatical sentences.
Through this two-stage process, concept selection and language model-based filtering, the data provider can strike a balance between semantic distinctiveness and linguistic naturalness. For data providers who possess intimate knowledge of their datasets, intuitively constructed watermarks are often sufficient for effective copyright protection. The advanced pipeline serves as an optional extension to further enhance the performance of the AQUA method. We have omitted a detailed description of this pipeline from the main text to maintain simplicity and highlight the core effectiveness of our primary approach.
Appendix B Real-world Deployment
B.1 Reference Distribution Estimation
Section 5.3 demonstrated the effectiveness of the AQUA method. However, in a real-world deployment, obtaining the simultaneous mean and variance of a RAG service before and after watermark injection is often infeasible. Typically, we can only observe a single set of statistics ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) from a suspected RAG service. To address this, we propose a verification strategy based on predefined reference distributions.
We first characterize the reference distribution of a clean Multimodal RAG using its mean and variance ($\mu_{clean}$, $\sigma^{2}_{clean}$), and similarly for a watermarked system ($\mu_{wm}$, $\sigma^{2}_{wm}$). Subsequently, we perform Welch’s t-test to compare the observed statistics ($\hat{\mu}_{suspect}$, $\hat{s}^{2}_{suspect}$) against these two respective reference distributions.
The null hypotheses ($\mathcal{H}_{0}$) for the two hypothesis tests are defined as follows:
  • Suspect vs. Clean: $\mathcal{H}_{0}^{(1)}:\hat{\mu}_{suspect}<\mu_{clean}$
  • Suspect vs. Watermarked: $\mathcal{H}_{0}^{(2)}:\hat{\mu}_{suspect}>\mu_{wm}$
To mitigate the risk of false accusations, the significance level $\alpha$ is set to a stringent value (e.g., $3e^{-5}$ as in Jovanović et al. (2025)).
Procedure for Establishing Reference Distributions.
Since verification distributions vary across different datasets and watermark patterns, data providers should derive reference distributions specific to their own datasets. We outline the following general procedure:
  1. 1.
    Select a RAG System: Employ an arbitrary RAG system (either open-source or closed-source) as the testbed.
  2. 2.
    Clean Distribution: Use the non-watermarked dataset as the knowledge base. Following the watermark verification protocols in Sections 4.1 and 4.2, perform probe queries to obtain the VSR metrics, and calculate the mean ($\mu_{clean}$) and variance ($\sigma^{2}_{clean}$).
  3. 3.
    Watermarked Distribution: Replace the knowledge base with the watermarked dataset and repeat the same process to obtain ($\mu_{wm}$, $\sigma^{2}_{wm}$).
Cross-System and Cross-Dataset Validation.
To demonstrate the robustness of our verification mechanism across diverse settings, we conduct additional experiments using both MMQA and WebQA datasets with three different VLMs as the RAG backbone: LLaVA-NeXT, InternVL3, and Qwen2.5-VL-Instruct. As shown in Table 8, regardless of the specific RAG system or dataset employed, there remains a significant divergence between the clean and watermarked distributions, ensuring reliable verification across diverse environments.
Table 8: Reference distribution statistics ($\mu_{clean}$, $\sigma^{2}_{clean}$, $\mu_{wm}$, $\sigma^{2}_{wm}$) across different RAG systems and datasets. A clear separation between clean and watermarked distributions is consistently observed.
RAG System & Watermark Dataset $\mu_{clean}$ $\sigma^{2}_{clean}$ $\mu_{wm}$ $\sigma^{2}_{wm}$
LLaVA-NeXT + $\textbf{AQUA}_{acronym}$ MMQA 0.01 0.02 0.60 0.20
LLaVA-NeXT + $\textbf{AQUA}_{spatial}$ MMQA 0.20 0.20 0.55 0.25
InternVL3 + $\textbf{AQUA}_{acronym}$ MMQA 0.01 0.01 0.86 0.13
InternVL3 + $\textbf{AQUA}_{spatial}$ MMQA 0.12 0.18 0.75 0.21
Qwen2.5-VL-Instruct + $\textbf{AQUA}_{acronym}$ MMQA 0.00 0.01 0.98 0.02
Qwen2.5-VL-Instruct + $\textbf{AQUA}_{spatial}$ MMQA 0.07 0.13 0.97 0.09
LLaVA-NeXT + $\textbf{AQUA}_{acronym}$ WebQA 0.01 0.02 0.77 0.03
LLaVA-NeXT + $\textbf{AQUA}_{spatial}$ WebQA 0.04 0.11 0.86 0.15
InternVL3 + $\textbf{AQUA}_{acronym}$ WebQA 0.01 0.01 0.79 0.19
InternVL3 + $\textbf{AQUA}_{spatial}$ WebQA 0.12 0.18 0.76 0.23
Qwen2.5-VL-Instruct + $\textbf{AQUA}_{acronym}$ WebQA 0.00 0.01 0.96 0.05
Qwen2.5-VL-Instruct + $\textbf{AQUA}_{spatial}$ WebQA 0.07 0.13 0.89 0.10
$\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$ require distinct parameters to characterize their respective reference distributions, as these depend on the specific watermark construction and its triggering performance. Based on our extensive experiments, we provide the following conservative example reference distributions with intentionally broadened variances to accommodate potential system fluctuations in real-world deployments:
  • $\textbf{AQUA}_{acronym}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = ($0.005$, $0.02$); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = ($0.6$, $0.2$)
  • $\textbf{AQUA}_{spatial}$: ($\mu_{clean}$, $\sigma^{2}_{clean}$) = ($0.2$, $0.2$); ($\mu_{wm}$, $\sigma^{2}_{wm}$) = ($0.55$, $0.25$)
As evident from Table 8, the actual separation between the clean and watermarked distributions is consistently distinct across all tested models and datasets, with the empirical statistics often showing even stronger separation than these conservative estimates. This confirms that once the watermark signal is successfully detected in a RAG system, the resulting VSR value typically diverges significantly from that of any non-infringing system, ensuring robust verification regardless of the specific RAG system or dataset configuration.
B.2 Collision Avoidance in Multi-Data Provider Environments
In a large-scale Retrieval-as-a-Service (RaaS) ecosystem involving multi-data providers, a critical concern is the potential for watermark collision, where triggers from different providers might conflict. AQUA addresses this challenge through a two-tiered approach: inherent probabilistic mitigation and a proposed centralized management architecture.
Inherent Mitigation via Combinatorial Verification.
As detailed in the methodology (Section 4), AQUA ensures uniqueness by requiring the injection and verification of a set of watermark images (e.g., $N=10$) coupled with a rigorous hypothesis testing framework. Unlike single-instance watermarking, our design relies on the joint probability of simultaneous retrieval and successful triggering across multiple distinct samples. Consequently, even without coordination, the probability of an accidental collision occurring concurrently across an entire set of watermark images from different providers is statistically negligible in real-world scenarios.
Centralized Watermark Registry.
For scenarios demanding absolute certainty—such as massive-scale commercial platforms—we propose the implementation of a Centralized Watermark Registry. Instead of relying solely on probabilistic independence, this approach positions the platform as a coordinating authority to explicitly manage watermark distribution. Functioning as a clearinghouse, the registry would require data providers to submit their intended triggers (e.g., acronyms or spatial configurations) prior to injection, allowing the system to cross-reference existing entries and preclude potential overlaps. To further streamline this process, the platform could proactively assign orthogonal trigger sets or exclusive namespaces to authorized providers. By transitioning from a probabilistic model to a deterministic registry, this architecture effectively eliminates accidental collisions, thereby simplifying dispute resolution and standardizing the verification protocol across the ecosystem.
Appendix C More Experimental Results
C.1 More Results of Effectiveness of AQUA
Q: Why is Welch’s t-test the appropriate statistical method in this experimental setting?
1) The standard Student’s t-test requires the assumption of equal variances (homogeneity of variance) between the two groups being compared. This assumption is not met in our analysis. For the datasets before and after watermarking, a given probe query retrieves different sets of images. Since the RAG generator’s output is conditioned on this retrieved context, the resulting VSR scores for the two groups are expected to have unequal variances. Therefore, Welch’s t-test, which does not assume equal variances, is the appropriate statistical method for our comparison. 2) For each watermarked image, we conducted multiple detection trials using a set of similar yet distinct probe queries. This repeated experimentation ensures that the resulting data distribution meets the normality assumption required for Welch’s t-test. 3) To ensure the stability and robustness of watermark detection in a practical deployment, we inject multiple watermarked images for a single probe query. For example, if the retriever returns the top-5 results, we can inject 10 watermarked images. This guarantees that for a given probe query, all images retrieved from the watermarked dataset are watermarked, while all images retrieved from the original dataset are normal. This design satisfies the independence assumption of Welch’s t-test.
Two-proportion Z-test. While Welch’s t-test serves as a robust and powerful method for comparing the means of two independent groups, particularly when population variances are unequal, the two-proportion Z-test is an equally standard and widely applied statistical tool specifically tailored for comparing proportions. The rationale for employing the Z-test in our experimental setting is direct and compelling.
The two-proportion Z-test is the canonical statistical method for evaluating whether an observed difference between two such proportions is statistically significant. Our experimental design, which involves two independent groups—the watermarked (experimental) and non-watermarked (control) datasets—and a large number of trials, perfectly aligns with the underlying assumptions of this test. It provides a rigorous framework for rejecting the null hypothesis that the performance is equivalent. Accordingly, we applied the two-proportion Z-test to our experimental data to quantitatively validate the efficacy of our watermarking scheme. The results of this analysis are presented below:
Table 9: The table shows the p-values obtained from the Z-test.
Models Methods MMQA WebQA
LLaVA- NeXT Naive $0.45$ $0.91$
Opt. $1.06e^{-3}$ $7.42e^{-2}$
$\textbf{AQUA}_{acronym}$ $1.28e^{-273}$ $4.32e^{-173}$
$\textbf{AQUA}_{spatial}$ $5.04e^{-43}$ $3.89e^{-38}$
InternVL3 Naive $0.38$ $0.73$
Opt. $4.97e^{-3}$ $6.19e^{-3}$
$\textbf{AQUA}_{acronym}$ $2.89e^{-251}$ $3.91e^{-110}$
$\textbf{AQUA}_{spatial}$ $4.31e^{-48}$ $7.73e^{-36}$
Qwen-VL-Chat Naive $0.53$ $0.84$
Opt. 5.19e35.19e{-3} $9.33e^{-2}$
$\textbf{AQUA}_{acronym}$ $5.62e^{-127}$ $8.02e^{-86}$
$\textbf{AQUA}_{spatial}$ $2.07e^{-52}$ $7.11e^{-27}$
Qwen2.5-VL-Instruct Naive $0.32$ $0.81$
Opt. $2.01e^{-2}$ $9.33e^{-3}$
$\textbf{AQUA}_{acronym}$ $1.11e^{-175}$ $2.70e^{-133}$
$\textbf{AQUA}_{spatial}$ $3.43e^{-71}$ $5.19e^{-47}$
Effectiveness across Varying Parameter Scales.
To comprehensively evaluate the scalability and generalizability of AQUA, we extended our experiments to include models with varying parameter scales, ranging from 2B to 38B. This analysis aims to verify the method’s applicability across diverse deployment scenarios, from resource-constrained edge devices to high-performance server-grade systems.
We maintained the consistent experimental settings as detailed in Section 5.3. Given that the retrieval ranking is solely determined by the fixed retriever component, we focus our reporting on the $CGSR$, which reflects the generator’s ability to correctly decode the watermark. The results are presented in Table 10.
Table 10: Effectiveness of AQUA across different model scales. We report the CGSRCGSR (%) on the MMQA dataset.
Model $\textbf{AQUA}_{acronym}$ $\textbf{AQUA}_{spatial}$
InternVL3-2B $70.53$ $63.76$
Qwen3-VL-2B-Instruct $92.18$ $91.67$
InternVL3.5-38B $90.71$ $84.33$
Qwen3-VL-32B-Instruct $99.83$ $98.91$
The empirical results demonstrate the robustness of AQUA regardless of model size:
  • Low-resource Feasibility: The method remains effective even on lightweight models. Notably, the 2B-parameter Qwen3-VL achieves a CGSR exceeding $91\%$, outperforming several larger baselines. This indicates that AQUA is viable for deployment in low-resource environments.
  • Scalability: As the model parameter count increases to the 30B+ range, the watermark triggering performance consistently improves, with Qwen3-VL-32B achieving near-perfect detection rates (up to $99.83\%$). This confirms that the watermarking mechanism aligns well with the enhanced reasoning capabilities of large-scale foundation models.
C.2 More Results of Harmlessness of AQUA